A narrow beam of light with wavelengths from 450 nm to 700 nm is incident perpendicular to one face of a prism made of crown glass, for which the index of refraction ranges from n = 1.533 to n = 1.517 for those wavelengths. The light strikes the opposite side of the prism at an angle of 37.0 ∘. Part A What is the angular spread of the beam as it leaves the prism?
Answer #1

P1 = sin-1 (sin37 x 1.533) = 67.31 degree

Similarly,

P2 = sin-1 (sin37 x 1.517) = 65.92 degree

Spread = P1 - P2 = 1.39 degree

Comment in case any doubt please rate my answer.....